# MLM: Learning Multi-task Loco-Manipulation Whole-Body Control for Quadruped Robot with Arm

### v1

@@ Abstract
Whole-body loco-manipulation for quadruped robots with arm remains a challenging problem, particularly in achieving multi-task control. To address this, we propose MLM, a reinforcement learning framework driven by both real-world and simulation data. It enables a six-DoF robotic arm–equipped quadruped robot to perform whole-body loco-manipulation for multiple tasks autonomously or under human teleoperation. To address the problem of balancing multiple tasks during the learning of loco-manipulation, we introduce a trajectory library with an adaptive, curriculum-based sampling mechanism. This approach allows the policy to efficiently leverage real-world collected trajectories for learning multi-task loco-manipulation. To address deployment scenarios with only historical observations and to enhance the performance of policy execution across tasks with different spatial ranges, we propose a Trajectory-Velocity Prediction policy network. It predicts unobservable future trajectories and velocities. By leveraging extensive simulation data and curriculum-based rewards, our controller achieves whole-body behaviors in simulation and zero-shot transfer to real-world deployment. Ablation studies in simulation verify the necessity and effectiveness of our approach, while real-world experiments on the Go2 robot with an Airbot robotic arm demonstrate the policy’s good performance in multi-task execution.
Index Terms:
Reinforcement learning, legged robot, multi-task loco-manipulation, whole-body control.
@@ I Introduction
Quadruped robot motion control, especially through reinforcement learning (RL), has gained significant attention and development, enhancing their diverse and impressive capabilities [1, 2, 3, 4, 5, 6]. A six-degree-of-freedom (DoF) robotic arm can offer more possibilities for enabling quadruped robots to perform manipulations. By integrating the locomotion of a quadruped robot with the manipulation capabilities of a robotic arm, a broader range of tasks can be accomplished compared to using a quadruped robot alone [7, 8].
However, achieving multiple loco-manipulation tasks with a single policy presents a challenge for quadruped robots. Employing a single policy across multiple tasks can reduce control complexity by eliminating the need for task-specific policies. However, balancing multiple tasks remains a problem that has not been adequately addressed by existing methods. Consequently, an adaptive balancing mechanism is essential. The method for achieving adaptive balance requires further research. In simulation-driven RL paradigms, the virtual assets and rewards specific to different tasks often vary, and careful, personalized design typically leads to a significant workload and lacks generalizability. Furthermore, the gap of visual information required for manipulation is often difficult to address due to significant differences in texture, color, lighting, and other factors between simulation and the real world [9]. A feasible solution is to distill multiple tasks into multiple trajectories learning. This avoids incorporating observational data of the task objectives (such as visual information) into the control, instead replacing it with trajectories.
@@
The acquisition of trajectories can be achieved through various methods, including UMI-like approaches [10, 11], teleoperation [12, 13], exoskeleton-based systems [14]. It is important to note that we do not collect whole-body trajectory data for quadruped loco-manipulation, which distinguishes our approach from imitation learning. This is due to the challenges of species selection for data collection and the inability to obtain such data from real robots. Thus, we use the UMI-like approach to obtain only the end-effector 6D manipulation trajectories, a lightweight solution independent of the robots.
@@
Integrating real-world multi-task manipulation trajectories into simulation and leveraging the simulation to efficiently generate training data that interacts with the environment is a feasible approach. Furthermore, by abstracting multiple tasks into multiple trajectories and designing a unified task reward, it can guide the quadruped robot in generating coordinated whole-body behaviors under a single policy to accomplish the tasks. These multi-task data from the real world provide spatial trajectories that align with the dynamics of the physical environment and are distributed across the robot’s manipulation space, which is essential for learning whole-body behaviors and generalizing to unseen trajectories [15]. This contrasts with sampling random spatial curves in simulation, which lack real-world dynamics like velocity and acceleration.
@@
It is important to note that in deployment schemes that only have access to historical trajectories, such as teleoperation, future trajectory information cannot be used as observation. However, future information is crucial for the stability of whole-body execution tasks, as it indicates the direction of the next step in the action [16]. Thus, a trajectory prediction mechanism is necessary.
@@
To address the aforementioned challenges, we introduce MLM, a learning-based multi-task loco-manipulation method. We integrate Fast-UMI [11] to provide manipulation trajectories. In the simulation, we train the control policy and transfer it to the real world. Our contributions are summarized as follows:
@@
  • We propose a Trajectory-Velocity Prediction-based framework for training a multi-task loco-manipulation controller, driven by both real-world and simulation data, with trajectory interface deployment enabled through teleoperation and a diffusion policy.
@@
  • We introduce a multi-task real-world trajectory data library, where adaptive and curriculum-based sampling enables the robot to balance performance across tasks and learn multi-task manipualtion with a single policy network.
@@
  • We deploy a real quadruped robot with arm and validate the excellent performance of our method in whole-body loco-manipulation control across multiple tasks.
@@ II-A RL-based Legged Locomotion
The success of RL-based motion controllers for legged robots, compared to model-based approaches, stems from the following key advantages. RL trains control policies using large amounts of data collected from simulated robot-environment interactions, capturing the dynamics of these interactions while leveraging far more data than real-world training allows [1, 3, 17]. Furthermore, reward-driven policy networks provide an end-to-end mapping from states to joint actions, eliminating the need for costly optimization during inference [18, 4].
Building on the parallel dynamic simulator Isaac Gym, a quadruped robot was trained to walk within minutes [19]. An actuator network was employed to model the complex actuator dynamics of the ANYmal robot, thereby enabling sim-to-real transfer [20]. Extending this approach, privileged learning was introduced to train the ANYmal robot, successfully achieving a blind motion controller for complex terrains [1]. Additionally, a method named RMA was proposed to develop robust quadruped locomotion controllers [3]. To facilitate gait learning and improve adaptation to complex terrains, gait priors were introduced [21, 2, 22], allowing robots to leverage motion data from flat terrains and rapidly converge to the desired gait pattern. While these efforts have significantly advanced the locomotion capabilities of legged robots [23, 24, 25, 26], the development of their loco-manipulation abilities still requires further exploration.
@@ II-B Legged Loco-Manipulation
To endow the manipulation capabilities of legged robots, researchers have integrated the mobility of legged platforms with the manipulation abilities of robotic arms. For instance, in [27], footstep planning was addressed, and a hierarchical optimization-based controller was developed to enable a quadruped robot with a six-DoF robotic arm to perform manipulation. The study in [7] proposed a approach that searches for whole-body trajectories and determines contact schedules to solve mobile manipulation tasks within predefined environments. Nonetheless, these optimization-based approaches often require task-specific design, presenting significant challenges in integrating the execution of multiple tasks within a single control policy. Additionally, they incur substantial computational overhead due to the optimization processes involved.
In recent years, several studies have focused on employing RL to achieve end-to-end control of legged robots with robotic arms and exploring the execution of policies across multiple tasks. In [28], the arm’s motion is modeled as predictable external torques that the RL-trained locomotion policy uses to compensate for the manipulator’s influence. Nevertheless, this non-whole-body approach limits the ability of the arm to utilize the mobility and posture adjustment capabilities of the quadruped robot to expand its manipulation range. To achieve whole-body control, an RL-based Regularized Online Adaptation method was proposed, successfully achieving sim-to-real transfer for quadruped loco-manipulation [29]. Despite this, the manipulation space defined under the arm base are susceptible to the influence of body shaking and are not stable enough. We utilize world space to address this issue, as the operational objectives are independent of body movement. In [30], a nonlinear reward fusion module was proposed to achieve whole-body behaviors for a wheeled quadrupedal manipulator. However, it did not integrate automatic execution of the task. In [31], an RL-based approach that integrates visual information for autonomous loco-manipulation execution was proposed. Yet, the inverse kinematics (IK) strategy is also susceptible to locomotion interference due to the arm base being affected by the body’s movement. In [16], a vision-driven, trajectory-based loco-manipulation in the task space was realized for a quadruped robot. However, it did not consider the integration of multiple tasks and the balance of performance across tasks within a single policy. Additionally, it is only applicable to trajectory generation methods that involve future trajectory reasoning, such as diffusion policy (DP). Teleoperation that only involve historical trajectories are not suitable.
@@ III Whole-body Control Method
Figure 2: Overview of the training and inference pipeline. We use reinforcement learning to train a whole-body control policy. A trajectory library, together with an adaptive sampling curriculum, supplies a multiple set of real-world manipulation trajectories for the policy to track. A normalized autoencoder encodes historical target trajectories, implicitly conveying information about upcoming targets; alternatively, the policy can receive future trajectories directly, enabling two deployment modes: teleoperation and DP generation. We also train an estimator to estimate base linear velocity $\hat{v}_{t}^{p}$ using supervised learning.
@@ III-A Whole-body Loco-Manipulation Policy
We eliminate the requirement for object and environment perception in simulation, which speeds up training and avoids the large sim-to-real gap caused by visual perception. Since information is partially observable, we model the loco-manipulation task as a partially observable Markov decision process (POMDP) [17, 22]. We employ RL to train a leg-arm whole-body policy $\pi_{\theta}$, aiming to find optimal policy parameters $\theta$ that maximize the discounted expected return $\mathbb{E}_{\mathbf{\tau}(\pi_{\theta})}[\sum_{t=0}^{T}{\gamma^{t}r_{t}}]$, where $r_{t}$ is the reward at time $t$, and $\gamma$ is the discount factor ($\gamma=0.99$).
We introduce an asymmetric Actor-Critic (AAC) framework to train the loco-manipulation policy. This approach is designed to address the POMDP problem while also avoiding the complexity of the two-stage training process [4].
State Space: The full state for the policy includes the quadruped robot’s proprioception $\boldsymbol{o}_{t}^{q}\in{\mathbb{R}^{42}}$, the six-DoF robotic arm’s proprioception $\boldsymbol{o}_{t}^{a}\in{\mathbb{R}^{18}}$, the trajectory pose sequence $o_{t-3:t+4}^{tcp}\in{\mathbb{R}^{72}}$, the privileged state $\boldsymbol{s}_{t}^{p}\in{\mathbb{R}^{21}}$, and the terrain elevation data $\boldsymbol{s}_{t}^{e}\in{\mathbb{R}^{187}}$.
@@
Specifically, $\boldsymbol{o}_{t}^{q}$ includes the quadruped’s body angular velocity $\boldsymbol{\omega}_{t}^{q}\in{\mathbb{R}^{3}}$, the gravity vector $\boldsymbol{g}_{t}^{q}\in{\mathbb{R}^{3}}$ representing its orientation, leg joint positions $\boldsymbol{\theta}_{t}^{q}\in{\mathbb{R}^{12}}$ and velocities $\dot{\boldsymbol{\theta}}_{t}^{q}\in{\mathbb{R}^{12}}$, and the previous time step’s leg joint actions $\boldsymbol{a}_{t-1}^{q}\in{\mathbb{R}^{12}}$. The arm proprioception $\boldsymbol{o}_{t}^{a}$ contains the arm joint positions $\boldsymbol{\theta}_{t}^{a}\in{\mathbb{R}^{6}}$, velocities $\dot{\boldsymbol{\theta}}_{t}^{a}\in{\mathbb{R}^{6}}$, and last arm joint actions $\boldsymbol{a}_{t-1}^{a}\in{\mathbb{R}^{6}}$. The trajectory pose sequence $o_{t-3:t+4}^{tcp}$ comprises Tool Center Points (TCP) over past and subsequent steps relative to time $t$, each including a 3D position and the first six elements of the rotation matrix. The privileged state $\boldsymbol{s}_{t}^{p}$, covering both robot and environment, includes the quadruped’s body linear velocity $\boldsymbol{v}_{t}^{q}\in{\mathbb{R}^{3}}$, contact forces $\boldsymbol{f}^{c}_{t}\in{\mathbb{R}^{12}}$, external force $\boldsymbol{f}^{e}_{t}\in{\mathbb{R}^{3}}$, and its position $\boldsymbol{p}^{e}_{t}\in{\mathbb{R}^{3}}$ on the robot. Finally, $\boldsymbol{s}_{t}^{e}\in{\mathbb{R}^{187}}$ represents the heights of 187 terrain points evenly distributed around the robot’s coordinate frame.
@@
Action Space: The whole-body policy outputs 12 joint position offsets for the legs, $\boldsymbol{a}_{t}^{q}\in{\mathbb{R}^{12}}$, and 6 joint position offsets for the arm, $\boldsymbol{a}_{t}^{a}\in{\mathbb{R}^{6}}$, both relative to their default positions. These offsets are added to the default positions and sent to the motors as target position commands for proportional-derivative (PD) control [30].
@@
Reward: Maintaining coordinated loco-manipulation of quadruped robots is challenging. Building on [22], we use adversarial motion prior (AMP) to encourage quadruped gaits learning via style rewards. During policy training, an adversarial discriminator network is trained concurrently to provide style rewards, defined as $r^{s}=\max[0,1-0.25\times\left(d^{\rm score}-1\right)^{2}]$.
@@
where $N$ is the total number of tasks. $\oplus$’ represents the concatenation operation. We have established a curriculum to update different tasks reward factor $\sigma_{n}^{p}$ and $\sigma_{n}^{o}$, more details are described in III-C. The constraint reward, $r^{l}$, is used to regulate the robot’s reasonable actions, energy limitations, and hardware constraints. All rewards are shown in Table I.
@@
Terrain Curriculum: In addition to the enhanced ability to adjust its posture with more degrees of freedom, the leg configuration of quadruped robots enables flexible adaptation to complex terrains. To enable a quadruped robot to perform loco-manipulation tasks in complex terrains, we introduce terrain curriculum. This approach allows the robot to progressively learn whole-body behaviors across several types of terrain, starting from easier to more difficult conditions.
These terrains include randomly undulating surfaces, slopes, discrete obstacles, stairs. Unlike pure locomotion tasks, we have removed the terrain traversal update mechanism, and instead, we use reward threshold updates. Specifically, we monitor the TCP tracking reward within an episode, and when it reaches the threshold, the robot progresses to the next level of the curriculum in the subsequent episode; otherwise, it regresses to the previous level. We have found that this mechanism effectively ensures the robot maintains a good tracking performance while preserving the stability of the base.
Domain Randomization: To reduce the sim-to-real gap, we randomize robot hardware, control, and environment parameters within specified ranges during training to ensure the simulation covers or approximates real-world conditions. As shown in Table II. This includes PD gains, link mass, payload mass and position, ground friction, and motor strength. Additionally, an observation delay of 0 to 0.02 seconds is introduced to mitigate the system’s non-real-time behavior.

### v2

@@ Abstract
Whole-body loco-manipulation for quadruped robots with arms remains a challenging problem, particularly in achieving multi-task control. To address this, we propose MLM, a reinforcement learning framework driven by both real-world and simulation data. It enables a six-DoF robotic arm–equipped quadruped robot to perform whole-body loco-manipulation for multiple tasks autonomously or under human teleoperation. To address the problem of balancing multiple tasks during the learning of loco-manipulation, we introduce a trajectory library with an adaptive, curriculum-based sampling mechanism. This approach allows the policy to efficiently leverage real-world collected trajectories for learning multi-task loco-manipulation. To address deployment scenarios with only historical observations and to enhance the performance of policy execution across tasks with different spatial ranges, we propose a Trajectory-Velocity Prediction policy network. It predicts unobservable future trajectories and velocities. By leveraging extensive simulation data and curriculum-based rewards, our controller achieves whole-body behaviors in simulation and zero-shot transfer to real-world deployment. Ablation studies in simulation verify the necessity and effectiveness of our approach, while real-world experiments on a Go2 robot with an Airbot robotic arm demonstrate the policy’s good performance in multi-task execution.
@@ I Introduction
Quadruped robot motion control, especially through reinforcement learning (RL), has gained significant attention and development, enhancing their diverse and impressive capabilities [lee2020learning, wjzamp, kumar2021rma, ji2022concurrent, Li2024Learning, Peng2025Learning]. A six-degree-of-freedom (DoF) robotic arm can offer more possibilities for enabling quadruped robots to perform manipulations. By integrating the locomotion of a quadruped robot with the manipulation capabilities of a robotic arm, a broader range of tasks can be accomplished [Sleiman2023Versatile, Yokoyama2024ASC].
@@
Employing a single policy can reduce control complexity by eliminating the need for task-specific policies. However, achieving multiple loco-manipulation tasks with a single policy presents a challenge for quadruped robots, and balancing multiple tasks remains unresolved in existing methods, making an adaptive mechanism essential and warranting further research. Moreover, in simulation-driven paradigms, task-specific assets and rewards demand heavy manual design with poor generalizability. The visual gap for manipulation also remains challenging due to differences in texture, color, lighting, and other factors between simulation and reality [Mohammad2024SplatSim]. A feasible solution is to distill multiple tasks into multiple trajectories learning. This avoids incorporating visual data of the task objectives into the control.
@@
Building on this idea, real-world multi-task trajectories can be introduced into simulation, while simulation can generate large-scale interaction data. By abstracting tasks into trajectories and defining a unified reward, quadruped robots can be guided to develop coordinated whole-body behaviors under a single RL policy. Real-world multi-task data provide spatial trajectories consistent with physical dynamics and distributed across the robot’s work space, which are essential for learning whole-body behaviors and generalizing to unseen trajectories [He2025ASAP]. In contrast, random spatial curves sampled in simulation lack realistic dynamics such as velocity and acceleration.
The acquisition of trajectories can be achieved through various methods, including UMI-like approaches [chi2024umi, Zhaxizhuoma2025FastUMI], teleoperation [Tony2023Learning, Anthony2023RT1], exoskeleton-based systems [Fang2024AirExo]. It is important to note that we do not collect the whole-body trajectory data, which distinguishes our approach from imitation learning. This is due to the challenges of species selection for data collection and the inability to obtain such data from real quadruped robot with an arm. Thus, we use the UMI-like approach to obtain only the end-effector 6D manipulation trajectories, a lightweight solution independent of the robots.
@@
In deployment schemes that only have access to historical trajectories, such as teleoperation, future trajectory information cannot be used as observation. However, future information is crucial for the stability of whole-body execution tasks, as it indicates the direction of the next step in the action [Huy2024UMIonLegs]. Thus, a trajectory prediction mechanism is necessary.
@@
To address the aforementioned challenges, we incorporate Fast-UMI [Zhaxizhuoma2025FastUMI] to provide real-world manipulation trajectories and introduce MLM, a learning-based multi-task loco-manipulation framework. The control policy is trained in simulation and subsequently transferred to the real world. Our contributions are summarized as follows:
@@
  • We propose a Trajectory-Velocity Prediction-based RL framework for training a multi-task loco-manipulation controller, driven by both real-world and simulation data, with trajectory interface deployment enabled through teleoperation and a diffusion policy.
@@
  • We introduce a multi-task real-world trajectory data library, where adaptive and curriculum-based sampling enables the robot to balance performance across tasks and learn multi-task manipulation with a single policy network.
@@
  • We deploy a real quadruped robot with an arm and validate the excellent performance of our method in whole-body loco-manipulation control across multiple tasks.
@@ II-A RL-based Legged Locomotion
The success of RL-based motion controllers for legged robots, compared to model-based methods, lies in two main advantages. First, RL trains control policies using large-scale simulated robot–environment interactions, capturing robot dynamics while leveraging far more data than real-world training allows [lee2020learning, kumar2021rma, miki2022learning]. Second, reward-driven policies enable end-to-end mapping from states to joint action, eliminating the need for costly optimization during inference [margolis2022rapid, ji2022concurrent].
Built on the parallel simulator Isaac Gym, quadruped robots can learn to walk within minutes [rudin2022learning]. To enable sim-to-real transfer, an network was used to capture the complex dynamics of the ANYmal robot’s motors [hwangbo2019learning]. This approach was later extended with privileged learning, yielding a blind locomotion controller for challenging terrains [lee2020learning]. Further advances include RMA, which enhances robustness in quadruped locomotion [kumar2021rma]. To accelerate gait learning and adapt to complex terrains, researchers introduced gait priors [escontrela2022adversarial, wjzamp, Liu2024Skill], leveraging flat-terrain motion data for faster convergence. These efforts have greatly advanced legged locomotion [Fan2024Rethinking, David2024ANYmalparkour, Miki2024Learning3D, Tairan2024Agile], yet the development of loco-manipulation remains largely unexplored.
@@ II-B Legged Loco-Manipulation
To enhance legged robots with manipulation skills, researchers have combined their mobility with the manipulation capabilities of robotic arms. For instance, in [Bellicoso2019ALMA], footstep planning was addressed, and a hierarchical optimization-based controller was developed to enable a quadruped robot with a six-DoF robotic arm to perform manipulation. The study in [Sleiman2023Versatile] proposed a approach that searches for whole-body trajectories and determines contact schedules to solve mobile manipulation tasks within predefined environments. Nonetheless, these optimization-based approaches often require task-specific design, presenting significant challenges in integrating the execution of multiple tasks within a single control policy. Additionally, they incur substantial computational overhead due to the optimization processes involved.
In recent years, several studies have focused on employing RL to achieve end-to-end control of legged robots with robotic arms. In [Ma2022Combining], the arm’s motion is modeled as predictable external torques, which the RL-trained locomotion policy compensates for. However, this non-whole-body approach limits the arm from leveraging the quadruped’s mobility and posture adjustments to extend its manipulation range. To enable whole-body control, an RL-based regularized online adaptation method was proposed, successfully achieving sim-to-real transfer for quadruped loco-manipulation [Fu2023DeepWholeBody]. Despite this, the arm-base-defined manipulation space is vulnerable to body shaking and thus unstable. To address this, we adopt a world frame, where operational objectives remain independent of body motion. In [Jiang2025Learning], a nonlinear reward fusion module was proposed to achieve whole-body behaviors for a wheeled quadrupedal manipulator. However, it did not integrate automatic execution of the task. In [Liu2024VisualWholeBody], an RL-based approach that integrates visual information for autonomous loco-manipulation execution was proposed. Yet, the inverse kinematics (IK) strategy is also susceptible to locomotion interference due to the arm base being affected by the body’s movement. In [Huy2024UMIonLegs], a vision-driven, trajectory-based loco-manipulation in the task space was realized for a quadruped robot. However, it did not consider the integration of multiple tasks and the balance of performance across tasks within a single policy. Additionally, it is only applicable to trajectory generation methods that involve future trajectory reasoning, such as diffusion policy (DP). Teleoperation that only involve historical trajectories are not suitable.
@@ III Whole-body Control Method
Figure 2: Overview of the pipeline. We use RL to train a whole-body control policy. A trajectory library, together with an adaptive sampling curriculum, supplies multiple real-world manipulation trajectories for the policy to track. An NAE encodes historical trajectories to estimate future targets, while the policy can also take future trajectories directly, supporting two deployment modes: teleoperation and automatic DP. In addition, we train a supervised estimator to predict the base linear velocity $\hat{v}_{t}^{p}$.
@@ III-A Whole-body Loco-Manipulation Policy
We eliminate the requirement for object and environment perception in simulation, which accelerates training and avoids the large sim-to-real visual gap. Since information is partially observable, the loco-manipulation task is modeled as a partially observable Markov decision process (POMDP) [miki2022learning]. RL is employed to train a whole-body policy $\pi_{\theta}$, aiming to find optimal parameters $\theta$ that maximize the discounted expected return $\mathbb{E}_{\mathbf{\tau}(\pi_{\theta})}[\sum_{t=0}^{T}{\gamma^{t}r_{t}}]$, where $r_{t}$ is the reward at time $t$, and $\gamma$ is the discount factor ($\gamma=0.99$). We introduce an asymmetric actor-critic framework to train the policy [ji2022concurrent]. This approach is designed to address the POMDP problem while also avoiding the complexity of the two-stage training.
State space: The input of the actor (policy network) includes a five-frame history sequence of the quadruped robot’s proprioception $\boldsymbol{o}_{t}^{q}\in{\mathbb{R}^{30}}$, the six-DoF robotic arm’s proprioception $\boldsymbol{o}_{t}^{a}\in{\mathbb{R}^{12}}$, the last action $\boldsymbol{a}_{t-1}\in{\mathbb{R}^{18}}$, and the end-effector pose trajectory $o_{t-3:t+4}^{tcp}\in{\mathbb{R}^{72}}$. Besides the actor’s input, the critic has access to the privileged state $\boldsymbol{s}_{t}^{p}\in{\mathbb{R}^{21}}$ and the terrain elevation data $\boldsymbol{s}_{t}^{e}\in{\mathbb{R}^{187}}$, which are only available during training and not used by the actor. A normalized autoencoder (NAE) only receives the historical end-effector pose sequence to predict future trajectory, while the memory encoder only receives the historical robot and arm proprioception. Further details are provided in Section III-B.
@@
Specifically, $\boldsymbol{o}_{t}^{q}$ includes the quadruped’s body angular velocity $\boldsymbol{\omega}_{t}^{q}\in{\mathbb{R}^{3}}$, the gravity vector $\boldsymbol{g}_{t}^{q}\in{\mathbb{R}^{3}}$ representing its orientation, leg joint positions $\boldsymbol{\theta}_{t}^{q}\in{\mathbb{R}^{12}}$ and velocities $\dot{\boldsymbol{\theta}}_{t}^{q}\in{\mathbb{R}^{12}}$. The arm proprioception $\boldsymbol{o}_{t}^{a}$ contains the arm joint positions $\boldsymbol{\theta}_{t}^{a}\in{\mathbb{R}^{6}}$ and velocities $\dot{\boldsymbol{\theta}}_{t}^{a}\in{\mathbb{R}^{6}}$. The last action $\boldsymbol{a}_{t-1}$ includes the previous time step’s leg joint action $\boldsymbol{a}_{t-1}^{q}\in{\mathbb{R}^{12}}$ and arm joint action $\boldsymbol{a}_{t-1}^{a}\in{\mathbb{R}^{6}}$. The trajectory pose sequence $o_{t-3:t+4}^{tcp}$ comprises Tool Center Points (TCP) over past and subsequent steps relative to time $t$, each including a 3D position and the first six elements of the rotation matrix. The privileged state $\boldsymbol{s}_{t}^{p}$, covering both robot and environment, includes the quadruped’s body linear velocity $\boldsymbol{v}_{t}^{q}\in{\mathbb{R}^{3}}$, contact force $\boldsymbol{f}^{c}_{t}\in{\mathbb{R}^{12}}$, external force $\boldsymbol{f}^{e}_{t}\in{\mathbb{R}^{3}}$, and its position $\boldsymbol{p}^{e}_{t}\in{\mathbb{R}^{3}}$ on the robot. Finally, $\boldsymbol{s}_{t}^{e}$ represents the heights of 187 terrain points evenly distributed around the robot frame.
@@
Action space: The whole-body policy outputs 12 joint position offsets for the legs, $\boldsymbol{a}_{t}^{q}\in{\mathbb{R}^{12}}$, and 6 joint position offsets for the arm, $\boldsymbol{a}_{t}^{a}\in{\mathbb{R}^{6}}$. These offsets are added to the default positions and sent to the motors as target position commands for proportional-derivative (PD) control [Jiang2025Learning].
@@
Reward: Maintaining coordinated whole-body behaviors is challenging. Building on [Liu2024Skill], we use adversarial motion prior (AMP) to encourage quadruped gaits learning via a style reward. During policy training, an adversarial discriminator is trained concurrently to provide the style reward, defined as $r^{s}=\max[0,1-0.25\times\left(d^{\rm score}-1\right)^{2}]$.
@@
where $N$ is the total number of tasks. $\oplus$’ represents the concatenation operation. We have established a curriculum to update different tasks reward factors $\sigma_{n}^{p}$ and $\sigma_{n}^{o}$, more details are described in III-C. The constraint reward, $r^{l}$, is used to regulate the robot’s reasonable action, energy limitations, and hardware constraints. All rewards are shown in Table I.
@@
Terrain curriculum: In addition to offering greater freedom of posture adjustment, the leg design of quadruped robots enables adaptation to terrains. To support loco-manipulation in such environments, we introduce a terrain curriculum [Liu2024Skill], which gradually trains the robot to master whole-body behaviors, progressing from simple to challenging terrains.
These terrains include randomly undulations, slopes, discrete obstacles, and stairs. Unlike locomotion tasks, we remove the terrain traversal update mechanism, and instead apply reward threshold updates. Specifically, we monitor the tracking reward within an episode, and when it reaches the threshold, the robot progresses to the next curriculum level in the subsequent episode; otherwise, it regresses to the previous one. This mechanism effectively maintains good tracking performance while ensuring base stability.
Domain randomization: To reduce the sim-to-real gap, we randomize robot hardware, control, and environment parameters within specified ranges during training [wjzamp]. As shown in Table II. This includes PD gains, link mass, payload mass and position, ground friction, and motor strength. Additionally, an observation delay of 0 to 0.02 seconds is introduced to mitigate the system’s non-real-time behavior.